Contribution Margin and Contribution Margin Ratio

For a recent year, McDugal's company-owned restaurants had the following sales and expenses (in millions):

Sales   $15,500
Food and packaging   $5,480
Payroll 4,100
Occupancy (rent, depreciation, etc.) 3,520
General, selling, and admin. expenses    2,400
Other expense 310
Total expenses (15,810)
Operating income (loss) $(310)

Assume that the variable costs consist of food and packaging, payroll, and 40% of the general, selling, and administrative expenses.

a. What is McDonald's contribution margin? Enter your answer in million, rounded to one decimal place.
$ million

b. What is McDonald's contribution margin ratio? Round your percentage answer to one decimal place.
%

c. How much would operating income increase if same-store sales increased by $900 million for the coming year, with no change in the contribution margin ratio or fixed costs?
$ million

d. What would have been the operating income or loss for the recent year if sales had been $900 million more?
$ million

e. To achieve break even for the recent year, by how much would sales need to increase? Enter your anwer in million rounded to the nearest whole number.
$ million
Answer #1

a) Variable cost = 5480+4100+(2400*40%) = 10540

Contribution margin = Sales-Variable cost = 15500-10540 = 4960

b) Contribution margin ratio = 4960/15500 = 32%

c) Net income increase by = 900*32% = 288

d) Net operating income = 288-310 = -22

e) Break even = 5270/.32 = 16468.75

Sales increase by (16468.75-15500) = 969